The Walt Disney Company (DIS) has four business segments, described as follows:

Media Networks: Television and radio
Parks and Resorts: Resorts, including Disneyland
Studio Entertainment: Motion pictures, musical recordings, and stage plays
Consumer Products & Interactive Media: Character merchandising, Disney stores, books, and games

Disney recently reported segment operating income, revenue, and invested assets (in millions) as follows:

    Operating
    Income
    
Revenue
    Invested
    Assets
Media Networks $7,755 $23,689 $32,706
Parks and Resorts 3,298 16,974 28,275
Studio Entertainment 2,703 9,441 15,359
Consumer Products & Interactive Media 1,965 5,528 9,332

a. Use the DuPont formula to determine the return on investment for the four Disney segments. Round Profit Margin and ROI to one decimal place and Investment Turnover to two decimal places.

Profit Margin Investment Turnover ROI
Media Networks % %
Parks and Resorts % %
Studio Entertainment % %
Consumer Products & Interactive Media % %

b. How do the four segments differ in their profit margin, investment turnover, and return on investment?
_________ has the highest profit margin, while______ has the lowest profit margin. ______ has the highest investment turnover, while _______ has the lowest investment turnover. ______ has the highest return on investment, while _____ has the lowest return on investment.
Answer #1

Profit Margin = operating income/sales revenue

Investment turnover = sales revenue/average Invested assets

ROI= Profit Margin*Investment Turnover

Media Networks

Profit Margin = 7755/23689= 32.7%

Investment turnover = 23689/32706 = 0.72 times

ROI= 23.5%

Parks and resorts

Profit margin = 3298/16974 = 19.4%

Investment turnover = 16974/28275 = 0.60 times

ROI = 11.6%

Studio entertainment

Profit margin = 2703/9441= 28.6%

Investment turnover = 9441/15359 = 0.61 times

ROI = 17.4%

Consumer Products

Profit margin = 1965/5528 = 35.5%

Investment turnover = 5528/9332 = 0.59 times

ROI = 20.9%

Consumer products has highest product margin

Parks and resorts has lowest

Media networks has highest turnover

Consumer products have lowest

Media has highest ROI

Parks has lowest